43-3d=d+9: 43-3d=d+9
Answer by orca(336)

Add 3d to both sides,
43-3d+3d=d+9+3d
43=4d+9
Subtract 9 from both sides,
43-9 = 4d+9-9
34 = 4d
Divide both sides by 4,
34/4 = 4d/4
17/2 = d
So d = 17/2